# MimicKit: A Reinforcement Learning Framework for Motion Imitation and Control

### v1

@@ 4. Methods
DeepMimic (Peng et al., 2018):
DeepMimic is a simple RL-based motion tracking method, which trains a tracking controller to follow target reference motions. This method is very general and reliable, and has been successfully applied to train controllers for a wide range of behaviors. DeepMimic is often a good starting point before considering more sophisticated techniques, and can be a highly effective method for applications that require precise replication of a target reference motion. However, a key limitation of DeepMimic is that the motion tracking objective used during training often leads to inflexible policies that are restricted to closely following a given reference motion. This can limit the agent’s ability to modify and adapt behaviors in the dataset as necessary to perform new tasks.
@@
Adversarial Motion Priors (Peng et al., 2021)
: Unlike DeepMimic, which trains a controllers to closely track a given reference motion, AMP is an adversarial distribution-matching method that aims to imitate the overall behavioral distribution (i.e. style) depicted in a dataset of motion clips, without explicitly tracking any specific motion clip. AMP provides more versatility than tracking-based methods, providing the agent with more flexibility to compose and adapt behaviors in the dataset in order to perform new tasks. However, a key drawback of distribution-matching methods, such as AMP, is that they are more prone to converging to local optima, especially for challenging, highly dynamics motions. Therefore AMP may struggle more to closely replicate challenging behaviors, compared to tracking-based methods, such as DeepMimic.
@@
Adversarial Skill Embeddings (Peng et al., 2022)
: ASE is an adversarial methods for training reusable generative controllers. This method combines adversarial imitation learning with a mutual information-based skill discovery objective to learn latent skill embeddings. Points in the latent space can be mapped to diverse behaviors by the ASE controller. Once trained, the ASE controller can be reused to perform new tasks by training task-specific high-level controllers to select skills from the learned latent space. Example arguments for training ASE models are provided in args/ase_humanoid_args.txt.
Adversarial Differential Discriminator (Zhang et al., 2025): ADD is an adversarial motion tracking method that uses a differential discriminator to automatically learn adaptive motion tracking objectives. This method can mitigate the manual effort required to design and tune tracking reward functions for different characters and motions. Example arguments for training ADD models are provided in args/add_humanoid_args.txt.

### v2

@@ 4. Methods
4.1. DeepMimic
DeepMimic is a simple RL-based motion tracking method (Peng et al., 2018), which trains a tracking controller to follow target reference motions. This method is very general and reliable, and has been successfully applied to train controllers for a wide range of behaviors. DeepMimic is often a good starting point before considering more sophisticated techniques, and can be a highly effective method for applications that require precise replication of a target reference motion. However, a key limitation of DeepMimic is that the motion tracking objective used during training often leads to inflexible policies that are restricted to closely following a given reference motion. This can limit the agent’s ability to modify and adapt behaviors in the dataset as necessary to perform new tasks.
@@ 4.1. DeepMimic
4.2. Adversarial Motion Priors (AMP)
Unlike DeepMimic, which trains a controllers to closely track a given reference motion, AMP is an adversarial distribution-matching method that aims to imitate the overall behavioral distribution (i.e. style) depicted in a dataset of motion clips (Peng et al., 2021), without explicitly tracking any specific motion clip. AMP provides more versatility than tracking-based methods, providing the agent with more flexibility to compose and adapt behaviors in the dataset in order to perform new tasks. However, a key drawback of distribution-matching methods, such as AMP, is that they are more prone to converging to local optima, especially for challenging, highly dynamics motions. Therefore AMP may struggle more to closely replicate challenging behaviors, compared to tracking-based methods, such as DeepMimic.
@@ 4.2. Adversarial Motion Priors (AMP)
4.3. Adversarial Skill Embeddings (ASE)
ASE is an adversarial methods for training reusable generative controllers (Peng et al., 2022). This method combines adversarial imitation learning with a mutual information-based skill discovery objective to learn latent skill embeddings. Points in the latent space can be mapped to diverse behaviors by the ASE controller. Once trained, the ASE controller can be reused to perform new tasks by training task-specific high-level controllers to select skills from the learned latent space. Example arguments for training ASE models are provided in args/ase_humanoid_args.txt.
4.4. Adversarial Differential Discriminator (ADD)
ADD is an adversarial motion tracking method that uses a differential discriminator to automatically learn adaptive motion tracking objectives (Zhang et al., 2025). This method can mitigate the manual effort required to design and tune tracking reward functions for different characters and motions. Example arguments for training ADD models are provided in args/add_humanoid_args.txt.